Find the distance between points A(4,-3) and B(-4,3).
I am not sure where to begin.

Answer by checkley71(8403)

THE X COORDINATES ARE 4-(-4)=8 UNITS THE Y COORDINATES ARE 3-(-3)=6 UNITSTHUS THE THIRD SIDE OR DISTANCE BETWEEN THESE TWO POINTS IS FOUND BYTHE FORMULA
A^2+B^2=C^2 OR 6^2+8^2=C^2 OR 36+64=C^2 OR C^2=100 OR C=10 UNITS BETWEEN THE TWO POINTS (4,-3)(-4,3)